OpenStudy (anonymous):

Let f be the function defined by f(x) = 3x^5-5x^3+2 a) On what intervals is f increasing b) On what intervals is the graph of f concave upward. c) Write the equation of each horizontal tangent line to the graph of f. Please provide some detail to your response.

OpenStudy (earthcitizen):

firstly, we want to know the the stationary points

OpenStudy (earthcitizen):

d/dx(3 x^5-5 x^3+2) = 15 x^2 (x^2-1)

OpenStudy (earthcitizen):

@ stationary point \[d/dx(3 x^5-5 x^3+2) = 15 x^2 (x^2-1)=0\]

OpenStudy (earthcitizen):

\[\therefore x=0 and x=1\]

OpenStudy (earthcitizen):

\[f(0)=2 and f(1)=0\]

OpenStudy (earthcitizen):

a. on what interval is f increasing ?

OpenStudy (mr.math):

a) f is increasing on the intervals at which the derivative is negative: \[f(x)=3x^5-5x^3+2 \implies f'(x)=15x^4-15x^2=15x^2(x^2-1).\] Do you know how to study signs of f'(x)?

OpenStudy (earthcitizen):

\[x =\pm 1\]

OpenStudy (mr.math):

It's pretty easy. \(15x^2\) is always positive, while x^2-1 is negative on (-1,1) and positive elsewhere. Thus f(x) is increasing on \((-\infty,\infty)-(-1,1)\).

OpenStudy (earthcitizen):

a. it increases at -1

OpenStudy (mr.math):

And decreasing on \([-1,1]\).

OpenStudy (earthcitizen):

http://www.wolframalpha.com/input/?i=+3x%5E5-5x%5E3%2B2 check this out!

OpenStudy (earthcitizen):

that's for b

OpenStudy (earthcitizen):

now for c

OpenStudy (earthcitizen):

this might tricky b/coz of the asymptote

OpenStudy (earthcitizen):

at x=0 , y=2

OpenStudy (earthcitizen):

this can be done by graph method or equation method which do you prefer ?

OpenStudy (anonymous):

sorry i haven't been responding probably eq'n method

OpenStudy (earthcitizen):

alryt, i think we shud draw the line at y=2

OpenStudy (earthcitizen):

it shud be a str8 line

OpenStudy (earthcitizen):

in c. we're asked to find the equation of the tangent line

OpenStudy (earthcitizen):

we know our max=-1and min=1

OpenStudy (earthcitizen):

plug these into the first equation

OpenStudy (anonymous):

A plot is attached.